Juan is 8 years older that his sister. In 3 years, he will be twice as old as she will be then. How old is each now?
Answer by Earlsdon(6294)

Let J = Juan's present age and S = his sister's present age.
1) J = S+8 "Juan is 8 years older than his sister"
2) J+3 = 2(S+3) "In 3 years, he will be twice as old as his sister will be then"
Substitute equation 1) into equation 2) and solve for S.
(S+8)+3 = 2(S+3) Simplify.
S+11 = 2S+6 Subtract S from both sides.
11 = S+6 Subtract 6 from both sides.
5 = S or S = 5 so...
J = S+8 so J = 5+8
J = 13 Juan is 13 years old and his sister is 5 years old.